10. Consider the defendant example from the lecture. If 3 jurors are added and each juror

has a 70% chance of finding the defended innocent, what is the probability that the

defendant is found innocent and set free?
Answer #1

Given that ,

p = 0.70

1 - p = 1 - 0.70 = 0.30

n =3

Using binomial probability formula ,

P(X = x) = (n C x) * px * (1 - p)n - x

P(X =0 ) = (3 C 0) * 0700 * (0.30)3

probability =0.0270